OpenStudy (anonymous):

I will give a medal and fan, please help :) Plans for a bridge are drawn on a coordinate grid. One girder of the bridge lies on the line y = –9x – 9. A perpendicular brace passes through the point (–1, 7). Write an equation of the line that contains the brace. y – 1 = –(x + 7) x – 7 = –9(y + 1) y – 7 = (x + 1) y – 7 = –9(x + 1)

OpenStudy (anonymous):

ok well we know the slope of the first line correct

OpenStudy (anonymous):

Thats the y = -9x - 9?

OpenStudy (anonymous):

Or is it the -1,7?

OpenStudy (anonymous):

they really stretch for these word problems, don't they?

OpenStudy (anonymous):

y=mx+b where m is the slope

OpenStudy (anonymous):

slope of \(y=mx+b\) is \(m\) what is the slope of \(y=-9x-9\) ?

OpenStudy (anonymous):

usually lol

OpenStudy (anonymous):

Yes they do @satellite73 I think thats what confuses me so much.

OpenStudy (anonymous):

So the slope is -9.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct, now we need to find the inverse

OpenStudy (anonymous):

now how about a perpendicular line. what slope will that have?

OpenStudy (anonymous):

What is the inverse? And Im not sure what a perpindicular line would have.

OpenStudy (anonymous):

if the slope of one line is \(m\) then the slope of a perpendicular line will be \(-\frac{1}{m}\) in english "negative reciprocal"

OpenStudy (anonymous):

so the negative reciprocal of \(-9\) is ...

OpenStudy (anonymous):

-1 over -9

OpenStudy (anonymous):

lol @satellite73 dont you have enough medals, and fans?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so simplified would get you what?

OpenStudy (anonymous):

actually, i don't see the answer listed above. maybe i am misinterpreting the question

OpenStudy (anonymous):

@NickR do you see the correct answer in the list?

OpenStudy (anonymous):

Idk. I hate math :(

OpenStudy (anonymous):

now that i look at it no i dont

OpenStudy (anonymous):

looks like all answers are supposed to be to the question "write the equation of the line parallel to \(y=-9x-9\) through the point \((-1,7)\)

OpenStudy (anonymous):

we all do

OpenStudy (amistre64):

might be a typo

OpenStudy (anonymous):

@amistre hi am i missing something?

OpenStudy (anonymous):

besides my hair i mean

OpenStudy (anonymous):

well either way a is out can tell you that

OpenStudy (anonymous):

WAIT. It didnt copy and paste right. The first answer says * -1 over -9 and 3rd answer says 1 over 9

OpenStudy (amistre64):

\[y -7= \frac19(x+1)\] \[9(y -7)= x+1\]

OpenStudy (anonymous):

go with C, it is almost always C (all but a set of measure zero)

OpenStudy (anonymous):

working through this i got y-7 = 1/9 (x+1)

OpenStudy (anonymous):

That means its c. Because after I corrected my typo thats what c is?

OpenStudy (anonymous):

ok glad we could help

OpenStudy (anonymous):

> It didnt copy and paste right.

OpenStudy (anonymous):

Thank you everyone, seriously. And sorry about my typo!

OpenStudy (anonymous):

all good not your fault